Find the angle between the vectors
α
Ì

=
2
i
+
2
j

k
αÌ…=2i+2j−k and
\
e
t
a
=
6
i

3
j
+
2
k
\eta=6i−3j+2k.
1
Expert's answer
2020-03-26T14:36:17-0400

Find the angle between the vectors  "\bar{\alpha}=2i+2j-k" and "\bar{\beta}=6i-3j+2k"

The cosine of the angle "\theta" between two vectors is equal to the dot product of this vectors divided by the product of vector magnitude


"\cos \theta={\bar{\alpha}\cdot\bar{\beta} \over |\bar{\alpha}|\cdot|\bar{\beta}|}"

"\bar{\alpha}=(2, 2, -1), \bar{\beta}=(6, -3, 2)"


"\bar{\alpha}\cdot\bar{\beta}=2(6)+2(-3)+(-1)(2)=4"

"|\bar{\alpha}|=\sqrt{(2)^2+(2)^2+(-1)^2}=3"

"|\bar{\beta}|=\sqrt{(6)^2+(-3)^2+(2)^2}=7"

Hence


"\cos \theta={4 \over 3\cdot7}={4 \over 21}"

"\theta=\arccos{{4 \over 21}}\approx79.02\degree"